# DRL-ORA: Distributional Reinforcement Learning with Online Risk Adaption

### v4

@@ A Generalized Non-convex Learning Perspective
where $Y(s_{t-1},\,a_{t-1})$ is the distribution generated by the ensemble networks. This distribution is supported by the $Q$-functions of each network, where each $Q$-value is computed as the quantile expectation over the return distribution (aleatoric uncertainty). It is important to note that the return distribution in each network is updated based on the output of its corresponding ensemble network. This ensures that the epistemic uncertainty is properly quantified and disentangled from the overall uncertainties.
We finally wrap-up the proposed algorithm – Distributional RL with Online Risk Adaption (DRL-ORA) in Algorithm 1. In each period $t$, the epistemic uncertainty is estimated using ensemble networks, where the return distribution is updated based on the output of each ensemble network (Line 9). The historical epistemic uncertainties are then used to dynamically update the risk parameter through the proposed online non-convex learning method (Line 10), where Algorithm 2 will be introduced later. The composite risk framework quantifies the combined effects of aleatory and epistemic risks by applying separate risk-awareness levels to the epistemic uncertainty and the return distribution (which implicitly captures aleatory uncertainty) using distortion functions $g_{\alpha}$ and $\beta$, respectively (Line 5 & 7). The action is chosen to minimize the epistemic uncertainty risk under the latest estimation of the parameter (Line 5).
@@
1:Input: Distortion function $g_{\alpha}$ with risk parameter $\alpha\in[\alpha_{\min},\,\alpha_{\max}]\subset[0,\,1]$ (the risk parameter is associated with each state-action pair) and corresponding risk measure $\rho_{\alpha}$; distortion function $\beta$, hyper-parameters: $M>0$, $K>0$, discounted factor $\gamma>0$; Training period $T>0$.
@@
10:  Compute the risk parameter $\alpha_{t+1}(s,\,a)$ through Algorithm 2, using the information $\alpha_{t}(s,\,a)$ and $X_{1}(s,\,a),...,X_{t+1}(s,\,a)$ for all (s,a)(s,\,a).
11:end for
@@
Given that our loss function is not necessarily convex in the risk parameter $\alpha$, standard online convex optimization algorithms cannot guarantee sublinear regret. The EWAF algorithm also fails in this setting, as our goal is to identify a specific risk parameter through continuous optimization rather than updating logit weights over a predefined discrete set. The Follow the Perturbed Leader (FTPL) algorithm (Agarwal, Gonen, and Hazan 2019) achieves sublinear regret complexity $O(T^{1/2})$ for the expected regret $\mathbb{E}[\mathfrak{R}_{T}]$ when the perturbation $\{\sigma_{t}\}_{t\geq 0}$ is sampled from an exponential distribution with parameter $\eta$. Theorem 4 demonstrates that a proper discretization of $\mathcal{A}$ into a finite set $\mathcal{A}^{\prime}\subseteq\mathcal{A}$, combined with grid search over $\mathcal{A}^{\prime}$, can efficiently solve the offline oracle problem while preserving the $O(T^{1/2})$ expected regret complexity. This method is particularly suitable when gradient or subgradient information is difficult to obtain.
Algorithm 2 Follow the Perturbed Leader (FTPL)
1:Input Parameter of exponential distribution $\eta$, arbitrarily chosen $\alpha_{1}\in\mathcal{A}^{\prime}$
2:for $t=2$ to $T$ do
3:  Generate randomly $\sigma_{t}\overset{\mathrm{i.i.d}}{\sim}\exp{(\eta)}$;
4:  Compute $\alpha_{t}\in\arg\min_{\alpha\in\mathcal{A}^{\prime}}\left\{\sum_{i=1}^{t-1}l_{i}(\alpha)-\sigma_{t}\alpha\right\}$.
5:  Observe loss function $l_{t}$.
6:end for
@@
Now we define a recursive loss function $l^{\prime}_{t}:\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}$ for all $t=1,\ldots,T$, where the first argument represents the decision variable $\alpha$, and the second argument is the parameter corresponding to the solution from the previous period. Specifically, at period t1t-1, we define $l^{\prime}_{t-1}(\alpha,\,\alpha_{t-1}):=|\rho_{\alpha}(X_{t-1})-\rho_{\alpha_{t-1}}(X_{t})|$. It can be shown that the cumulative regret under the above recursive loss function is equivalent to the cumulative regret under the regular loss function; that is,
@@
We then modify and implement Algorithm 2 with the offline oracle model (Step 4) as
@@
The connection between this algorithm’s offline oracle model and the satisficing measure (Brown and Sim 2009) from the decision analysis literature, can be established. The satisficing measure evaluates the extent to which a random outcome fails to meet a fixed target. To formalize this connection. Let $\tau_{t}:=\rho_{\alpha_{t-1}}(X_{t})$ denote the “target” at period $t$. Given $\Omega$ as the scenario space, if $\min_{\omega\in\Omega}X_{t-1}(\omega)\geq\tau_{t}$, then $\alpha_{t}=\alpha_{\min}$ (which could be 0, depending on the chosen risk measure). Conversely, if $\max_{\omega\in\Omega}X_{t-1}(\omega)\leq\tau_{t}$, then $\alpha_{t}=\alpha_{\max}$ (which could be $1$, depending on the chosen risk measure). Otherwise, the offline oracle problem of Algorithm 2 becomes equivalent to solving the following quasi-concave optimization problem:
@@
where the optimal solution provides the level of non-attainability of $X_{t-1}$ relative to $\tau_{t}$. Theorem 5, based on the analysis in Section 3.4 of (Brown and Sim 2009) and Theorem 3.6 of (Huang 2020), demonstrates that when $\rho_{\alpha}$ is chosen as CVaR, Problem (4) can be reformulated as a stochastic convex program.
@@
Suppose $\rho_{\alpha}=\text{CVaR}_{\alpha}$. We have $\alpha_{t}=\min_{b\geq 0}\mathbb{E}\left[\left(b(X_{t-1}-\tau_{t})+1\right)_{+}\right]$ with $\tau_{t}=\rho_{\alpha_{t-1}}(X_{t})$.
After applying sample average approximation for the expectation, Problem (4) can be further reformulated as a linear programming (LP) problem. The number of decision variables and constraints are both $O(K)$, and the best-known LP algorithm has a complexity of $O(K^{2.5})$ (Vaidya 1989). However, we develop a specialized search algorithm, illustrated in Appendix Section B, which can determine $\alpha_{t}$ with significantly lower computational complexity.
@@
Experiments and Applications
Figure 2: ORA with 90% confidence interval.
We conduct experiments on three classes of tasks to evaluate the performance of DRL-ORA. In all experiments, CVaR is used in ORA-based methods with $\alpha_{\min}=0.1$, resulting in $\mathcal{A}=[0.1,1]$. Additionally, we set $\beta(q)=q$, making the agent risk-neutral with respect to the implicit aleatory uncertainties of the tasks. We also implement the two existing risk aware adaption approaches: ART (Liu, van Kampen, and De Croon 2023) and TOP (Moskovitz et al. 2021; Wang, Hu, and Zhang 2024; Wang et al. 2025) as benchmarks.
@@
We first demonstrate the performance of DRL-ORA on Cartpole and the Atari games: Hero, MsPacman and SpaceInvaders. Here we show the results of CartPole task in addition to Example 1. Figure 2 shows that DRL-ORA outperforms all other methods, with a noticeable reward advantage at the early stages of the episodes. Furthermore, the 90% confidence interval of DRL-ORA highlights its stability. Due to limited spaces, please refer to Appendix Section C for the additional experiment results on other Atari games.
@@
We evaluate DRL-ORA on an open-source Nano Drone navigation task (Liu, van Kampen, and De Croon 2023), where ART has shown superior performance compared to IQN. In this partially observable environment, the drone aims to maximize its score by reaching the target while avoiding obstacles. It gains points for success but loses points for time and collisions, and the agent’s level of risk-awareness directly influences its exploration strategy and overall performance. We record the complete training and testing results for ORA with the regular loss function (“Recursive ORA” in short), ORA with the recursive loss function, ART, TOP and IQN under various fixed risk parameters. All training parameters, including learning rate and random seeds, are consistent with the default settings in (Liu, van Kampen, and De Croon 2023), see Appendix Table 4.
@@
ORA 92 6 81 13 64 28
As shown in Figure 3, ORA shows the best training performance and the fastest convergence among other benchmarks. Table 1 also presents the average testing results of each algorithm over 500 evaluation runs in environments with varying obstacle densities (i.e., the number of obstacles). Testing is particularly important in this navigation task, as excessive exploration during training may result in lower rewards but ultimately lead to an effective policy with strong testing performance. The testing results show that ORA outperforms all other algorithms, especially in environments with higher uncertainty (Density 6 & 12). In contrast, in the task with lower environmental uncertainty (Density 2), Recursive ORA slightly outperforms ORA while offering more efficient computation and requiring less storage for historical epistemic uncertainties.
@@
We assess DRL-ORA by applying it to a classic operations research (OR) problem, the Knapsack problem, on OR-gym (Hubbs et al. 2020; Balaji et al. 2019), a widely utilized RL library for OR problems. The Knapsack problem is a combinatorial optimization problem aiming at maximizing the total value of items placed in a knapsack without exceeding its maximum capacity. The agent’s reward is the value of items placed in the knapsack, and an episode terminates once the capacity is full. There is no aleatory uncertainty in the Knapsack problem. We set 50 items in the Binary Knapsack (Hubbs et al. 2020) for a lower computational load while keeping other settings aligned with the default. All training hyper-parameters are provided in Appendix Table 2. In terms of testing results in Figure 4, ORA’s average reward obtained is higher than all IQN’s. Note that the testing result should be normally higher than the final training reward because a fixed $\epsilon$ probability for random action selection is used in the whole training, while there is no random action selection in testing.
Figure 4: Testing results with 90% confidence interval. ”Comp.” means Composite IQN. The “Optimal episodic reward” is the benchmark solved via dynamic programming.
@@
In Figure 5, IQN with $\alpha=1.0$ demonstrates a clear advantage over IQN with $\alpha=0.1$, while regular Deep $Q$-Network (DQN) (Mnih et al. 2015). fails to perform effectively. This suggests that lower risk awareness is more appropriate in tasks without aleatory uncertainty. Once again, ORA outperforms the best IQN, TOP and ART, particularly during the early and intermediate stages of training. Additionally, we conduct an ablation experiment comparing ORA with the IQN Composite method (Hai et al. 2022). The IQN Composite algorithm is derived from Algorithm 1 by fixing $\alpha_{t}=1$ for all $t=1,\ldots,T$ and for all state-action pairs, making it risk-neutral with respect to epistemic uncertainty. This method replaces the expectation imposed on the return distribution with an expectation imposed solely on epistemic uncertainties, as there is no aleatory uncertainty in this task. Figure 5 shows that IQN Composite fails to achieve similar performance to ORA in the early training episodes, indicating that ORA’s superior performance stems from its adaptive mechanism for risk level selection.
@@
Appendix Table 1 reports, for ORA, the monitored left-truncated variance (LTV) with respect to the median of the epistemic uncertainty distribution obtained in each training period of the three experiments. The table also shows the corresponding risk parameters used in each case. CartPole has the smallest LTV value in average. Its LTV fluctuates over the whole training process, and the corresponding risk parameter remains at a high level. We also observe that LTV is almost uni-modal over episodes in Nano Drone navigation and Knapsack. A high risk parameter (lower risk-awareness level), is favored at the beginning to acquire knowledge of the environment, whereas it subsequently declines in performance when large LTVs are detected. Knapsack has the largest LTV value which decreases slower than that in Nano Drone navigation after reaching the peak. This can explain why the superior performance of ORA lasts longer in the Knapsack problem than that in Nano Drone navigation (see Figure 3 and 5).
Conclusion and Future Research
In this work, we propose an RL framework with an adaptive mechanism for learning risk levels to hedge against the epistemic uncertainties. The method only needs an extra multi-head ensemble structure so is eligible of employing different loss functions and distortion risk measures and being applied to IQN-type algorithms. Various applications validate the superiority of the proposed method.
For future research, we plan to develop more accurate and efficient methods for epistemic uncertainty quantification, e.g., improving the scalability when a large number of ensemble networks are applied. We will also extend our approach to RL tasks under non-stationary environments.
Appendix A Appendix
Preliminaries on the IQN Loss Function

### v5

@@ A Generalized Non-convex Learning Perspective
where $Y(s_{t-1},\,a_{t-1})$ is the distribution generated by the ensemble networks. This distribution is supported by the $Q$-functions of each network, where each $Q$-value is computed as the quantile expectation over the return distribution (aleatoric uncertainty). It is important to note that the return distribution in each network is updated based on the output of its corresponding ensemble network. This ensures that the epistemic uncertainty is properly quantified and disentangled from the overall uncertainties. Appendix Figure 6 shows the tests on different ensemble size $K$. It demonstrates that the increase of ensemble network size does not exhibit a significant correlation with performance improvement. Both the time and memory overhead grow slightly faster than linear growth with $K$. With the default value of $K$, this overhead is not significant.
We finally wrap-up the proposed algorithm – Distributional RL with Online Epistemic Risk Adaptation (DRL-ORA) in Algorithm 1. In each period $t$, the epistemic uncertainty is estimated using ensemble networks, where the return distribution is updated based on the output of each ensemble network (Line 9). The historical epistemic uncertainties are then used to dynamically update the risk parameter through the proposed online non-convex learning method (Line 10-11). The composite risk framework quantifies the combined effects of aleatory and epistemic risks by applying separate risk-awareness levels to the epistemic uncertainty and the return distribution (which implicitly captures aleatory uncertainty) using distortion functions $g_{\alpha}$ and $\beta$, respectively (Line 5 & 7). The action is chosen to minimize the epistemic uncertainty risk under the latest estimation of the parameter (Line 5).
@@
1:Input: Aleatory uncertainty distortion function $\beta$; epistemic uncertainty distortion function $g_{\alpha}$ with risk parameter $\alpha\in[\alpha_{\min},\,\alpha_{\max}]\subset[0,\,1]$ (the risk parameter is associated with each state-action pair) and corresponding risk measure $\rho_{\alpha}$; parameter of exponential distribution $\eta$, arbitrarily chosen $\alpha_{1}\in\mathcal{A}^{\prime}$, hyper-parameters: $M>0$, $K>0$, discounted factor $\gamma>0$; Training period $T>0$.
@@
10:  Generate randomly $\sigma_{t+1}\overset{\mathrm{i.i.d}}{\sim}\exp{(\eta)}$;
11:  Compute
$\alpha_{t+1}\in\arg\min_{\alpha\in\mathcal{A}^{\prime}}\left\{\sum_{i=1}^{t}l_{i}(\alpha)-\sigma_{t+1}\alpha\right\}.$
12:end for
@@
Given that our loss function is not necessarily convex in the risk parameter $\alpha$, standard online convex optimization algorithms cannot guarantee sublinear regret. The EWAF algorithm also fails in this setting, as our goal is to identify a specific risk parameter through continuous optimization rather than updating logit weights over a predefined discrete set. The Follow the Perturbed Leader (FTPL) algorithm [agarwal2019learning] achieves sublinear regret complexity $O(T^{1/2})$ for the expected regret $\mathbb{E}[\mathfrak{R}_{T}]$ when the perturbation $\{\sigma_{t}\}_{t\geq 0}$ is sampled from an exponential distribution with parameter $\eta$. Theorem 4 demonstrates that a proper discretization of $\mathcal{A}$ into a finite set $\mathcal{A}^{\prime}\subseteq\mathcal{A}$, combined with grid search over $\mathcal{A}^{\prime}$, can efficiently solve the offline oracle problem while preserving the $O(T^{1/2})$ expected regret complexity. This method is particularly suitable when gradient or subgradient information is difficult to obtain.
@@
Now we define a recursive loss function $l^{\prime}_{t}:\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}$ for all $t=1,\ldots,T$, where the first argument represents the decision variable $\alpha$, and the second argument is the parameter corresponding to the solution from the previous period. Specifically, at period t1t-1, we define
$l^{\prime}_{t-1}(\alpha,\,\alpha_{t-1}):=|\rho_{\alpha}(X_{t-1})-\rho_{\alpha_{t-1}}(X_{t})|.$
It can be shown that the cumulative regret under the above recursive loss function is equivalent to the cumulative regret under the regular loss function; that is,
@@
We then modify and implement Algorithm 1 with the offline oracle model (Line 11) as
@@
The connection between this algorithm’s offline oracle model and the satisficing measure [brown2009satisficing] from the decision analysis literature, can be established. The satisficing measure evaluates the extent to which a random outcome fails to meet a fixed target. To formalize this connection. Let $\tau_{t}:=\rho_{\alpha_{t-1}}(X_{t})$ denote the “target” at period $t$. Given $\Omega$ as the scenario space, if $\min_{\omega\in\Omega}X_{t-1}(\omega)\geq\tau_{t}$, then $\alpha_{t}=\alpha_{\min}$ (which could be 0, depending on the chosen risk measure). Conversely, if $\max_{\omega\in\Omega}X_{t-1}(\omega)\leq\tau_{t}$, then $\alpha_{t}=\alpha_{\max}$ (which could be $1$, depending on the chosen risk measure). Otherwise, the offline oracle problem of Algorithm 1 becomes equivalent to solving the following quasi-concave optimization problem:
@@
where the optimal solution provides the level of non-attainability of $X_{t-1}$ relative to $\tau_{t}$. Theorem 5, based on the analysis in Section 3.4 of brown2009satisficing and Theorem 3.6 of huang2020data, demonstrates that when $\rho_{\alpha}$ is chosen as CVaR, Problem (4) can be reformulated as a stochastic convex program.
@@
Suppose $\rho_{\alpha}=\text{CVaR}_{\alpha}$. We have
$\alpha_{t}=\min_{b\geq 0}\mathbb{E}\left[\left(b(X_{t-1}-\tau_{t})+1\right)_{+}\right],$
with $\tau_{t}=\rho_{\alpha_{t-1}}(X_{t})$.
After applying sample average approximation for the expectation, Problem (4) can be further reformulated as a linear programming (LP) problem. The number of decision variables and constraints are both $O(K)$, and the best-known LP algorithm has a complexity of $O(K^{2.5})$ [vaidya1989speeding]. However, we develop a specialized search algorithm, illustrated in Appendix Section B, which can determine $\alpha_{t}$ with significantly lower computational complexity.
@@
5 Experiments and Applications
We conduct experiments on three classes of tasks to evaluate the performance of DRL-ORA. In all experiments, CVaR is used in ORA-based methods with $\alpha_{\min}=0.1$, resulting in $\mathcal{A}=[0.1,1]$. Additionally, we set $\beta(q)=q$, making the agent risk-neutral with respect to the implicit aleatory uncertainties of the tasks. We also implement the two existing risk aware adaptation approaches: ART [liu2022adaptive] and TOP [moskovitz2021tactical, wang2024improving, wang2025adaptive] as benchmarks.
@@ 5 Experiments and Applications
We first demonstrate the performance of DRL-ORA on CartPole and the Atari games: Hero, MsPacman and SpaceInvaders. Here we analyze the results of CartPole task in addition to Example 1. Figure 7 in Appendix, built upon Figure 1, shows that DRL-ORA outperforms all other methods, with a noticeable reward advantage at the early stages of the episodes. To statistically validate the observed performance differences, we employed the Mann-Whitney $U$ test [10.1214/aoms/1177730491]. The resulting Rank-Biserial correlation effect sizes (0.990 against ART and 0.787 against TOP, $p$ < 0.001) confirm a substantial advantage of ORA over both the ART and TOP baselines in the CartPole environment. Furthermore, the 90% confidence interval of DRL-ORA highlights its stability. To evaluate the robustness of our approach, we compared the ORA using quantile ("ORA(quantile)") against the default CVaR-based version ("ORA(CVaR)"). The results show that despite a slight performance reduction, ORA(quantile) maintains considerable risk adaptation capability. The subsequent experiments further demonstrate the robustness of ORA across different risk measures. Due to limited spaces, please refer to Appendix Section C for the additional experiment results on other Atari games.
@@
We evaluate DRL-ORA on an open-source Nano Drone navigation task [liu2022adaptive], where ART has shown superior performance compared to IQN. In this partially observable environment, the drone aims to maximize its score by reaching the target while avoiding obstacles. It gains points for success but loses points for time and collisions, and the agent’s level of risk-awareness directly influences its exploration strategy and overall performance. We record the complete training and testing results for ORA with the regular loss function (“Recursive ORA” in short), ORA with the recursive loss function, ART, TOP and IQN under various fixed risk parameters. All training parameters, including learning rate and random seeds, are consistent with the default settings in [liu2022adaptive], see Appendix Table 5.
@@
ORA(CVaR) 92 6 81 13 64 28
ORA(quantile) 94 6 75 16 52 23
As shown in Figure 2, ORA not only shows the best training performance and fastest convergence but also significantly outperforms both baselines ($p$ < 0.001), achieving Rank-Biserial effect sizes of 0.319 over ART and 0.309 over TOP. Table 1 also presents the average testing results of each algorithm over 500 evaluation runs in environments with varying obstacle densities (i.e., the number of obstacles). Testing is particularly important in this navigation task, as excessive exploration during training may result in lower rewards but ultimately lead to an effective policy with strong testing performance. The testing results show that ORA outperforms all other algorithms, especially in environments with higher uncertainty (Density 6 & 12). In contrast, in the task with lower environmental uncertainty (Density 2), Recursive ORA slightly outperforms ORA while offering more efficient computation and requiring less storage for historical epistemic uncertainties.
@@
We assess DRL-ORA by applying it to a classic operations research (OR) problem, the Knapsack problem, on OR-gym [hubbs2020or, balaji2019orl], a widely utilized RL library for OR problems. The Knapsack problem is a combinatorial optimization problem aiming at maximizing the total value of items placed in a knapsack without exceeding its maximum capacity. The agent’s reward is the value of items placed in the knapsack, and an episode terminates once the capacity is full. There is no aleatory uncertainty in the Knapsack problem. We set 50 items in the Binary Knapsack [hubbs2020or] for a lower computational load while keeping other settings aligned with the default. All training hyper-parameters are provided in Appendix Table 4. In terms of testing results in Figure 3, ORA achieves a higher average reward than IQN, demonstrating that by precisely capturing epistemic uncertainty, our adaptive algorithm maintains its effectiveness in such scenarios. Our ORA method achieves perfect separation from TOP (effect size = 1.000) and substantial advantages over DQN (0.699) and ART (0.544), all with $p$ < 0.001. Note that the testing result should be normally higher than the final training reward because a fixed $\epsilon$ probability for random action selection is used in the whole training, while there is no random action selection in testing.
Figure 3: Testing results with 90% confidence interval. "Comp." means Composite IQN. The “Optimal episodic reward” is the benchmark solved via DP.
@@
In Figure 4, IQN with $\alpha=1.0$ demonstrates a clear advantage over IQN with $\alpha=0.1$, while regular Deep $Q$-Network (DQN) [mnih2015human]. fails to perform effectively. This suggests that lower risk awareness is more appropriate in tasks without aleatory uncertainty. Once again, ORA outperforms the best IQN, TOP and ART, particularly during the early and intermediate stages of training. Additionally, we conduct an ablation experiment comparing ORA with the IQN Composite method [hai2022rasr]. The IQN Composite algorithm is derived from Algorithm 1 by fixing $\alpha_{t}=1$ for all $t=1,\ldots,T$ and for all state-action pairs, making it risk-neutral with respect to epistemic uncertainty. This method replaces the expectation imposed on the return distribution with an expectation imposed solely on epistemic uncertainties, as there is no aleatory uncertainty in this task. Figure 4 shows that IQN Composite fails to achieve similar performance to ORA in the early training episodes, indicating that ORA’s superior performance stems from its adaptive mechanism for risk level selection.
@@
Table 2 in Appendix, reports, for ORA, the monitored left-truncated variance (LTV) with respect to the median of the epistemic uncertainty distribution obtained in each training period of the three experiments. The table also shows the corresponding risk parameters used in each case. CartPole has the smallest LTV value in average. Its LTV fluctuates over the whole training process, and the corresponding risk parameter remains at a high level. We also observe that LTV is almost uni-modal over episodes in Nano Drone navigation and Knapsack. A high risk parameter (lower risk-awareness level), is favored at the beginning to acquire knowledge of the environment, whereas it subsequently declines in performance when large LTVs are detected. Knapsack has the largest LTV value which decreases slower than that in Nano Drone navigation after reaching the peak. This can explain why the superior performance of ORA lasts longer in the Knapsack problem than that in Nano Drone navigation (see Figure 2 and 4).
6 Conclusion
In this work, we propose an RL framework with an adaptive mechanism for learning risk levels to hedge against the epistemic uncertainties. The method only needs an extra multi-head ensemble structure so is eligible of employing different loss functions and distortion risk measures and being applied to IQN-type algorithms. Various applications validate the superiority of the proposed method. For future research, we plan to develop more accurate and efficient methods for epistemic uncertainty quantification, e.g., improving the scalability when a large number of ensemble networks are applied. We will also extend our approach to RL tasks under non-stationary environments. Our approach is expected to remain valid under non-stationary environment as it is adaptive according to the latest experience/episode (epistemic uncertainty estimation).
Appendix
DRL-ORA: Distributional Reinforcement Learning with Online Epistemic Risk Adaptation
Appendix A Preliminaries on the IQN Loss Function